### v1

# Understanding In-context Learning of Addition via Activation Subspaces

Abstract
To perform in-context learning, language models must extract signals from individual few-shot examples, aggregate these into a learned prediction rule, and then apply this rule to new examples. How is this implemented in the forward pass of modern transformer models? To study this, we consider a structured family of few-shot learning tasks for which the true prediction rule is to add an integer $k$ to the input. We find that Llama-3-8B attains high accuracy on this task for a range of $k$, and localize its few-shot ability to just three attention heads via a novel optimization approach. We further show the extracted signals lie in a six-dimensional subspace, where four of the dimensions track the unit digit and the other two dimensions track overall magnitude. We finally examine how these heads extract information from individual few-shot examples, identifying a self-correction mechanism in which mistakes from earlier examples are suppressed by later examples. Our results demonstrate how tracking low-dimensional subspaces across a forward pass can provide insight into fine-grained computational structures.
footnotetext: Equal advising.
1 Introduction
Large language models (LLMs) have demonstrated remarkable in-context learning (ICL) capabilities, where they can adapt to various new tasks through few-shot demonstrations in the prompt. Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; von Oswald et al., 2023; Bai et al., 2023; Zhang et al., 2023; Nichani et al., 2024) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024). However, little is known about the fine-grained computational structure of in-context learning for large models.
To address this, we perform a detailed analysis of ICL for a structured family of tasks: learning to add a constant number (“add-k”). For these tasks, models are given an input sequence “$x_{1}\rightarrow y_{1}\#...\#x_{n}\rightarrow y_{n}\#x_{q}\rightarrow$” where $y_{i}=x_{i}+k$ for all $i$. This task is mathematically well-defined and tractable for an in-depth mechanistic analysis, while exhibiting the core challenge of extracting and applying a pattern from ICL examples. Moreover, modern language models are empirically capable of solving it (they correctly predict $y_{q}=x_{q}+k$).
Figure 1: In this paper, we study the activation space of attention heads in Llama-3-8B during in-context learning (ICL) for addition tasks. Our main findings: (1) out of $1024$ attention heads in Llama-3-8B, only three are important for these tasks (§4); (2) each head encodes the task information in a six-dimensional subspace (§4.1); (3) a six-dimensional subspace further decomposes into a four-dimensional subspace capturing the unit digit of the task information and a two-dimensional subspace capturing its ten digit (§4.3); and (4) task information is mostly extracted from a subset of tokens (§5). The figure illustrates these findings using an example add-k prompt.
To understand the mechanism behind in-context add-k, we study it in Llama-3-8B. We first discover that only a small number of model components are responsible for add-k: out of 1024 attention heads, only three do most of the work (§3). We find the task information for each head (i.e., the number $k$) is encoded in a six-dimensional subspace; zooming into one of the heads, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4). We also examine how this head aggregates information from individual in-context examples (§5).
To discover the attention heads that are responsible for add-k, we introduce a novel approach inspired by previous work on function vectors (Todd et al., 2024), which are vectors that imitate ICL when they are patched into the residual stream (§2.2). We formulate an optimization method that identifies $33$ important heads whose outputs yield good ICL performance (§3.1). Through further ablation, we find that all but 3 of these 33 heads can be replaced with their mean value without significantly decreasing the performance (§3.2).
Since these three heads are responsible for computing add-k, they must extract signal about $k$ from individual examples and then represent $k$ in the residual stream. To understand how $k$ is represented, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1). Furthermore, zooming into the six-dimensional subspace of one head, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions (with frequencies $2,5,10$), and a two-dimensional subspace tracking the coarse-grained magnitude (the ten digit) (§4.2-4.3). This offers insight on how LLMs internally represent addition, aligning with recent findings showing that (non-ICL) addition is encoded via trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025).
Finally, we examine how this head aggregates information from individual in-context examples. We find that the information is mostly extracted from the $y_{i}$ token positions and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2). Moreover, we identify a self-correction mechanism: if the signal extracted from one token has an error, later tokens are likely to write in the opposite direction of the error (§5.3). This means that ICL goes beyond simple averaging of inputs and has stateful dynamics.
Our findings suggest that ICL tasks can be implemented through specialized activation subspaces from a small number of attention heads that extract, represent and aggregate task-related information in interpretable ways. We introduce a novel optimization method to identify important attention heads and demonstrate how a focused analysis of a simple ICL task can reveal delicate latent space structures and sophisticated computational strategies in a large language model.
2 Preliminaries
2.1 Model and Task
We begin by specifying the model and task studied in this paper.
Model. We study Llama3-8B-Instruct, which has 32 layers and 32 attention heads per layer, and a residual dimension of 4096. We denote each head as a tuple (layer index, head index), where both indices range from 0 to 31.
Task.
For a constant $k$, the add-k task is to add $k$ to a given input integer $x$ to predict $y=x+k$. In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$" (where each $y_{i}=x_{i}+k$). The examples are concatenated in the prompt using the separator “##\##", followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt. A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e. $y_{i}-x_{i}$ in the prompts) but not the input domain, in contrast to most previous studies (Todd et al., 2024; Hendel et al., 2023).
We construct data for the task by varying $x\in[1,100]$ and $k\in[1,30]$ (thus $y\in[2,130]$) since Llama3-8B is empirically capable of solving the addition task in this range. We consider the following three types of prompts:
  1. 1.
    five-shot ICL prompt, where all 5 demonstration examples satisfy $y_{i}=x_{i}+k$ for a fixed $k\in[1,30]$. We also call this add-k ICL.
  2. 2.
    mixed-$k$ ICL prompt, where the five demonstration examples are $y_{i}=x_{i}+k_{i}$ for five possibly different $k_{i}$ values.
  3. 3.
    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$" for some $x_{q}$.
2.2 Activation Patching and Function Vectors
Next, we briefly introduce activation patching, a common interpretability technique that will be used throughout the paper, and the function vector, a construction that we will use to identify important heads for our ICL tasks.
Activation patching.
Activation patching is performed by taking the activations of a model component when the model is run on one prompt, then patching in these activations when the model is run on a different prompt. Patching can either replace the model’s base activations or add to them; we will primarily consider the latter.
Specifically, if $z_{l}$ is the original value of the residual stream at layer $l$ at the final token position $\to$”, we patch in the replacement $z_{l}+v$, where $v$ is constructed from activations on a different input prompt. We choose the layer $l$ at one third of the network’s depth and construct $v$ from “function vector” heads, following Todd et al. (2024), as described next.
Function vectors.
Function vectors $v_{k}$ are vectors constructed from the outputs of selected attention heads, designed so that adding $v_{k}$ to the residual stream of a zero-shot prompt approximates the effect of the five-shot add-k task. For example, $v_{k}$ might be the average output of one or more attention heads across a set of five-shot add-$k$ examples. We define the intervention accuracy of $v_{k}$ as the average accuracy obtained on zero-shot prompts when $v_{k}$ is added to the residual stream across all tasks $k$. This metric captures how effectively $v_{k}$ encodes task-specific information about $k$. For comparison, we define clean accuracy as the accuracy on five-shot prompts without any intervention, also averaged across all tasks $k$.
In more detail, consider an attention head $h$: let $h(p)$ denote the output of head $h$ on prompt $p$ at the last token position, and define $h_{k}$ as the average of $h(p)$ across all five-shot add-k prompts. [^1] Similar to Todd et al. (2024), we approximate this average using $100$ randomly generated five-shot add-k prompts, where each prompt contains random demonstration examples $x_{i}$ and the final query input $x_{q}$ is chosen exactly once from each integer in $[1,100]$. Todd et al. (2024) identified a subset $\mathcal{H}$ of attention heads (for a different set of tasks) such that the vector $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ has high intervention accuracy—that is, adding it to the residual stream of zero-shot prompts effectively recovers few-shot task performance.
Building on this framework, we consider multiple ways to construct $v_{k}$: (1) the task-specific mean $h_{k}$ over the add-k task (as described above); (2) the overall mean $\bar{h}$ (i.e., average across all $k$), which removes task-specific information about $k$; or (3) the specific value $h(p)$ on an individual prompt $p$. Throughout the paper, we call $\bar{h}$ the mean-ablation of head $h$ and call $h_{k}$ the head vector of $h$. Beyond the three choices above, we sometimes project a head’s output onto a lower-dimensional subspace or scale it by a coefficient. Unlike Todd et al. (2024), we recover a different set of heads through solving a novel global optimization problem (§3.1) and systematic ablation studies (§3.2), which achieves higher performance (Figure 3).
Figure 2: The chain of localization in §3 and §4. We first identify $33$ significant attention heads (out of 1024) via a global optimization method (§3.1), then narrow down to 3 main heads while mean-ablating the remaining 30 heads (§3.2). We next study the structure of the representations of each main head by localizing it to a six-dimensional subspace (§4.1), and decompose it into a four-dimensional subspace encoding the unit digit and a two-dimensional subspace encoding the tens digit (§4.3).
3 Identifying Three Aggregator Heads
To understand the mechanism behind the add-k ICL task, we first need to find out what model components are responsible for performing it. In this section, we find that three heads do most of the work for add-k. We find these heads by first solving an optimization problem to identify $33$ significant heads out of 10241024102410243.1), and then further narrowing down to three main heads via systematic ablations (§3.2). This process is illustrated in the first row of Figure 2.
3.1 Identifying Significant Heads via Sparse Optimization
We will search for a set of heads that store the information for the add-k task, in the sense that their output activations yield good function vectors for add-k2.2).
Formulating the sparse optimization.
More formally, define $v_{k}(c)=\sum_{h}c_{h}\cdot h_{k}$, the sum of head outputs weighted by $c$, where $h$ goes over all $1024$ heads in the model. (Recall that $h_{k}$ is the average output of head $h$ averaged across a large dataset of five-shot add-k prompts.) We will search for a sparse coefficient vector $c\in[0,1]^{1024}$ such that adding $v_{k}(c)$ to the residual stream of a zero-shot prompt achieves high accuracy on the add-k task.
Let $\ell(x_{q},y_{q};v)$ be the cross-entropy loss between $y_{q}$ and the model output when intervening the vector $v$ onto the input “$x_{q}\to$”(i.e. replacing $z_{l}$ with $z_{l}+v$ in the forward pass on “$x_{q}\to$”, where $z_{l}$ is the layer-$l$ residual stream at the last token). We optimize $c$ with respect to the loss
$\mathcal{L}(c)=\mathbb{E}_{k\in[30]}\mathbb{E}_{x_{q}\in[100]}[\ell(x_{q},x_{q}+k;v_{k}(c))]+\lambda\|c\|_{1},$ (1)
where the regularization term with weight $\lambda$ promotes sparsity.
(a) Coefficients of heads.
(b) Average indirect effects of heads (Todd et al., 2024).
Figure 3: Comparison of significant heads identified by our optimized coefficients (left) and by average indirect effects (AIE) from the previous method (Todd et al., 2024) (right). Colors indicate the magnitude of each head’s importance (coefficients or AIE). The top $33$ heads identified by both methods are highlighted with frames ($13$ heads common across both methods in red and other $20$ heads in blue). Our identified heads yield an intervention accuracy of $0.85$, compared to the previous method’s accuracy of $0.31$. Both methods select heads from similar layers, but our optimization approach is significantly more effective.
Training details. We randomly select $25$ add-k tasks from our total $30$ tasks for training and in-distribution testing, and use the remaining five tasks only for out-of-distribution testing. For each task add-k, we generate $100$ zero-shot prompts “$x_{q}\rightarrow$”, where $x_{q}$ ranges over all integers from $[1,100]$ and the target output is $x_{q}+k$, yielding one data point for each (prompt, task) pair. We randomly split the data points of the $25$ tasks into training, validation, and test sets in proportions of $0.7$, $0.15$, and $0.15$, respectively. We use AdamW with learning rate $0.01$ and batch size $128$. We set the regularization rate $\lambda$ as $0.05$, which promotes sparsity while incurring little loss in accuracy. During training, we clip the coefficients $c$ back to $[0,1]$ if they go out of range after each gradient step.
Results. We get coefficients that achieves high intervention accuracy. In particular, at the final epoch, the average accuracies of intervening on test data for the $25$ tasks and the five tasks ($0.83$ and $0.87$) are both close to the average accuracies of doing no intervention ($0.89$ and $0.92$). To identify the important heads for the tasks, we plot the values of the coefficients in the final epoch for each layer and head index (Figure 3(a)). We find $33$ heads have coefficients $c_{h}$ greater than $0.2$, most ($21$) of which are one. In contrast, the other ($991$) coefficients are all smaller than $0.01$, most ($889$) of which are zero. We call the $33$ heads significant heads and denote the set of them as $\mathcal{H}_{\text{sig}}$.
Comparison to previous approach. We compare our optimization approach with the previous method from Todd et al. (2024) for identifying important heads, which selects heads based on average indirect effect (AIE). To perform a fair comparison, we construct our function vector by summing the outputs of our selected heads directly without weighting by coefficients, matching their methodology. Using this construction, we achieve an intervention accuracy of $0.85$, close to the clean accuracy of $0.87$, indicating that our $33$ heads captures most of the necessary information for the add-k task. In contrast, selecting the top 33 heads according to AIE yields a much lower intervention accuracy of $0.31$. [^2] Todd et al. select the top 10 heads in their work, but this yields an even lower accuracy of $0.05$ in our setting. In Figure 3, we visualize the coefficients and AIE values of heads from both methods and highlight the $13$ heads that are common to both approaches.
3.2 Further Refinement via Ablations
We suspect many heads are primarily responsible for storing formatting information (such as ensuring the output appears as a number) rather than encoding information about $k$ itself. Intuitively, while we require the overall signal transmitted by these heads, we do not need any information about the specific value of $k$. To test this hypothesis, we perform mean-ablations: replacing each task-specific signal $h_{k}$ with the overall mean $\overline{h}$ across all values of $k$2.2).
Specifically, we conduct mean-ablations over subsets of the 33 significant heads and measure the resulting intervention accuracy of the corresponding function vectors. Formally, when ablating a subset $\mathcal{H}_{0}$, the resulting function vector is given by
$v=\sum_{h\in\mathcal{H}_{0}}\overline{h}+\sum_{h\in\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}}h_{k}.$ (2)
Here, heads in $\mathcal{H}_{0}$ contribute only their overall mean signals, while heads in $\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}$ retain their task-specific information.
Focusing on two layers via layer-wise ablation. To efficiently narrow down the important heads, we first perform ablations at the level of layers. From Figure 3(a), we observe that the significant heads $\mathcal{H}_{\text{sig}}$ are concentrated primarily in the middle and late layers. We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query. After trying different sets of layers, we found that ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while ablating any other combination of layers causes negligible drops in accuracy (Table 3). After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
Identifying three final heads via head-ablation. To understand the individual contributions of each head within layers 13 and 15, we perform mean-ablations at the level of individual heads. We first assess the intervention accuracy when retaining only the output of a single head while mean-ablating all other significant heads; however, this generally results in low accuracy. We hypothesize that the output magnitude of a single head is too small to significantly influence the model output, even if it encodes task-relevant information. To amplify each head’s effect, we scale its output by a coefficient (e.g., 5). We find that three heads—head1 = $(15,2)$, head2 = $(15,1)$, and head3 = $(13,6)$—achieve intervention accuracies close to the clean accuracy when appropriately scaled, while all other heads show much lower accuracies regardless of scaling (Table 1). This suggests that these three heads individually encode the task information much better than any others.
Head  Coefficient   Accuracy
No intervention  N/A   0.87
$(15,2)\text{:=Head1}$ $6$   0.85
$(15,1)\text{:=Head2}$ $5$   0.83
$(13,6)\text{:=Head3}$ $5$   0.66
Any other head  Optimal   0.19
Table 1: Intervened accuracies for scaling up each single head’s output by an optimal coefficient. Each of the top three heads (in red) achieves much higher intervention accuracy compared to any other significant head in layer $13$ and $15$ (in blue).
Finally, to remove the need for scaling while maintaining high intervention accuracy, we sum the outputs of these three heads (each with a coefficient of one) and mean-ablate all others. We find that summing the top three, top two, and only the top head yields intervention accuracies of $0.79$, $0.61$, and $0.21$, respectively. The three heads are thus collectively sufficient for performing the add-k task.
Validating necessity of the three heads via ablating them in five-shot ICL. So far, we have studied these three heads mainly through their contribution to the function vector $v_{k}$. We next directly test their necessity in the original five-shot ICL setting, by ablating the outputs of these three heads when running the model on many random five-shot ICL prompts. Our experiment shows that mean-ablating these three heads in five-shot prompts yields an accuracy of $0.43$, sharply decreasing from the clean accuracy $0.87$ by half. For comparison, we mean-ablate $20$ random sets of three significant heads (other than head 1,2,3); their accuracies remain close to the clean accuracy: $95\%$ of them have accuracy at least $0.86$.
4 Characterizing the Aggregator Subspace
For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL examples. Our next goal is to understand how task information is represented in the activation space. Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
In this section, we dissect their activation spaces in three stages: (1) Localize a six–dimensional task subspace in each head via principal component analysis (PCA) (§4.1); (2) Rotate this subspace into orthogonal feature directions aligned with sinusoidal patterns across $k$4.2); (3) Decompose the six-dimensional space into a four-dimensional unit-digit subspace and a two-dimensional magnitude subspace that separately encode the units and tens of the answer (§4.3).
4.1 Localizing to Six-dimensional Subspace
To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Figure 4). We then check that the function vectors found earlier remain effective after projecting onto the subspace. Specifically, we replace each head vector $h_{k}$ with its projection onto the subspace $\tilde{h}_{k}$ to obtain a new function vector $\tilde{v}_{k}=\sum_{h\in\mathcal{H}}\tilde{h}_{k}$ (a variant of $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ in §2.2). We find that $\tilde{v}_{k}$ has intervention accuracy $0.76$, which is close to the intervention accuracy of $0.79$ before projection (§3.2). Thus, we confine our study to this concise six-dimensional subspace for each head. The rest of the section focuses on head 1 for simplicity.
Figure 4: The explained variance ratio vs. number of PCs for each head. The first six PCs make up most of the explained variance ($97\%$) for each head.
Figure 5: Coordinates of head1’s vectors (their inner products with a PC) for the first six PCs across different add-$k$ tasks. The coordinates in the first five PCs reveal periodic patterns.
4.2 Identifying Feature Directions Encoding Periodic Patterns
To understand how the six-dimensional subspace of a head represents the task information, we first examine the coordinates of the head vectors (their inner products with the PCs) as a function of $k$. This reveals partially periodic patterns in the first five components (Figure 5).
This motivates us to linearly transform the six PCs to find directions that encode pure periodic patterns. Mathematically, if we can find a linear transformation of the six PC-coordinate functions that fits trigonometric functions, then by applying the transformation on the PCs, we can obtain six directions whose coordinate functions encode the periodicity.
To find trigonometric functions to fit, we searched over different periods and phases and performed least squares regression. We found six trigonometric functions at periods $2$, $5$, $10$, $10$, $25$, and $50$ that could be expressed as functions of the top 6 PCs with low regression error (Figure 6). We apply the resulting linear transformation to the six PCs to obtain a new set of directions that encode these six pure periodic patterns, which we call feature directions.
Figure 6: The coordinate functions (inner product with feature directions) of head 1’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
4.3 Decomposing to Subspaces Encoding Subsignals
Leveraging the feature directions identified previously, we decompose the head activation subspace into lower-dimensional components that separately encode different subsignals relevant to the task—in this case, the units digit and tens digit.
By construction, each of the feature directions (viewed as linear projections of $h_{k}$) are periodic functions of $k$. Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo $T$”. Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We verify these hypotheses through causal intervention. In Figure 7, we validate hypothesis (ii): we establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal. We include analogous experimental results for the two other hypotheses in Appendix B.
Figure 7: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto the “unit subspace” (left) and out of the “unit subspace” (right). Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully accounts for the final-answer errors. This confirms that the unit subspace specifically encodes the unit-digit signal.
5 Signal Extractors of ICL Examples
Previously, we localized the model’s behavior to three heads and their corresponding six-dimensional subspaces, then examined how the model represents the task information ($k$ for add-k) inferred from the ICL examples in one subspace. Now, we analyze how the model extracts the task information from the ICL examples.
In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior. For simplicity, we focus on head 1 in this section.
5.1 Mathematical Observation: Tracing Subspace back to Previous Tokens
We begin with a mathematical observation that lets us trace the subspace at the final token back to corresponding subspaces at earlier token positions. Intuitively, a head’s output at the last token is a weighted sum of transformed residual streams from the previous tokens, with the weights given by the attention scores. Thus, the signal extracted from previous tokens is the transformed residual stream at that token.
Formally, a head $h$’s output at the last token of a prompt $p$ can be written as
$h(p)=\sum_{t\in p}\alpha_{t}\cdot O_{h}V_{h}\cdot z_{t},$ (3)
where $\alpha_{t}$ is the attention score from the last token to each token $t$, satisfying $\sum_{t\in p}\alpha_{t}=1$, $z_{t}$ is the residual stream input to the head $h$ at token $t$, $V_{h}$ is the value matrix, and $O_{h}$ is the output matrix mapping from head-dimensional space to model-dimensional space.
Let $W_{h}$ denote the projection matrix onto the six-dimensional subspace for head $h$. Then the projected signal at the final token, $W_{h}\cdot h(p)$, can be decomposed into contributions from previous tokens as $W_{h}\cdot\alpha_{t}O_{h}V_{h}z_{t}$, each lying in the transformed head subspace defined by the matrix $W_{h}\cdot O_{h}V_{h}$. In the following subsection, we analyze the magnitudes and directions of these signals, and study how signals from different examples interact.
5.2 Signal Extractor for each Example
Figure 8: For a mixed-$k$ ICL prompt, the strength of the task-signal contribution of each previous token $t$ to the final token, $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$. Decomposing it into two parts: (1) the norm of the extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$ (top), and (2) the attention score from the final token $\alpha_{t}$ (bottom), both parts consistently peak at the tokens $t=y_{i}$ (in bold).
To understand how examples contribute to model generation at the last token, we identify which tokens contribute the most, then examine what information they provide. By the analysis above, the task-signal contribution of each previous token to the final token through the head $h$ is $\alpha_{t}W_{h}O_{h}V_{h}z_{t}$. This can be decomposed into two parts: (1) extracted information: $W_{h}O_{h}V_{h}z_{t}$, the residual stream input projected into the relevant subspace; and (2) aggregation weight: $\alpha_{t}$, the attention score of the final token to the previous token. We plot the norms of the extracted information and the aggregation weights for a random mixed-$k$ ICL prompt in Figure 8. Both the strength of the extracted information and the aggregation weights peak at $y_{i}$ tokens.
We next examine what specific information is extracted from each of these tokens. To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},h_{k}\rangle$ between the extracted information and the head vector for each task $k$. In Figure 9, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$. We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector $h_{k}$ (projected onto the subspace and normalized to have unit norm). For each example $x_{i}\rightarrow y_{i}$, the inner product peaks for $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
5.3 Signal Correlation among Examples
Having studied the signal extracted from each individual example in the previous subsection, we next study how signals from different examples interact to execute the ICL task. To do so, we compute the correlation between the extracted signals from different demonstration examples: for each $y_{i}$ token, we first compute the inner product between the residual stream input to head 1, $z_{t}$, and the corresponding task vector $h_{k}$, where $k=y_{i}-x_{i}$. Then, we compute the correlation of these measures across each pair of five positions over $100$ add-k prompts, yielding $\binom{5}{2}$ correlation values per task.
To analyze the correlation, we sum the negative correlation values and positive correlation values respectively for each task, and calculate the various statistics (max, average and min) over all tasks (Table 2). The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, indicating that the signals from any two demonstration examples are mostly negatively correlated. This suggests a self-correction mechanism: intuitively, when the head extracts a noisy signal from one example, signals from subsequent examples are more likely to correct the error, thereby stabilizing the final representation.
Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlation
Average $-2.01$ $0.27$
Min $-1.40$ $0.07$
Max $-2.34$ $0.54$
Table 2: The statistics (min, max, and average) of the absolute values of negative correlation sum and positive correlation sum respectively over the $30$ tasks. The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, which indicates that the signals from any two demonstration examples are mostly negatively correlated, suggesting a self-correction mechanism.
6 Related Work
Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Elhage et al., 2021; Wu et al., 2023; Wang et al., 2022; Feng & Steinhardt, 2023; Gould et al., 2023).
Interpreting arithmetic tasks.
A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Nikankin et al., 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025). Zhou et al. (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers. Different from them, our work is the first to analyze addition for in-context learning in a large pretrained model, and we find similar representation structures to Zhou et al. (2024) and Kantamneni & Tegmark (2025).
Interpreting in-context learning.
Researchers have constructed detailed models of in-context learning (ICL) for small transformer models (Garg et al., 2022; von Oswald et al., 2023; Bai et al., 2023; Zhang et al., 2023; Nichani et al., 2024), as well as coarser-grained treatments for large pretrained models, attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023). Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads. Motivated by this, we perform a detailed analysis on the mechanisms behind a family of ICL addition tasks through a novel optimization method inspired by function vectors, which achieves better performance than the method in Todd et al. (2024). Unlike the tasks studied in Todd et al. (2024), our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.
Causal analysis.
Our methodological approach draws on recent developments in the causal analysis of neural networks that aims to understand the causal influence of model components to its behavior by performing targeted interventions on the internal states of models (Geiger et al., 2021; Vig et al., 2020). In particular, our methods draw inspiration from the causal mediation analysis used in Todd et al. (2024), activation patching (Meng et al., 2022), and causal scrubbing (Chan et al., 2022).
7 Discussion
In this work, we provided a detailed mechanistic analysis of in-context learning for addition tasks in Llama-3-8B. We found that a small number of attention heads operating in low-dimensional subspaces can extract, represent, and aggregate ICL task information in structured and interpretable ways. We performed our analysis in five steps:
  1. 1.
    Use sparse optimization to identify important attention heads whose outputs construct effective function vectors for ICL tasks (§3.1).
  2. 2.
    Localize task information to a smaller subset of heads via ablations (§3.2).
  3. 3.
    Further localize to low-dimensional subspaces via PCA on each remaining head (§4.1).
  4. 4.
    Examine this subspace qualitatively, which uncovered periodic patterns in the activation space (§4.2) that decomposed into interpretable subspaces encoding unit-digit and magnitude information (§4.3).
  5. 5.
    Exploit algebraic structure in the transformer to connect “aggregation” subspaces at the final token position with “extraction” subspaces at the earlier $y_{i}$ tokens (§5).
This same methodology (identify important heads, restrict to relevant subspaces, and examine the remaining information qualitatively) could be extended to other models and tasks. Most steps in our methodology also scale easily: the sparse optimization is fully automatic. While mean ablation involved some qualitative judgment, we can fold both of these steps into a single optimization task that mean ablates some heads while fully removing others. PCA is also automatic. For the final step that involves a qualitative examination of the subspaces, future work could explore automating this step using AI systems.
There are a few other questions stemming from our work. First, the three important heads are capable of individually representing and extracting information, but Llama3-8B relies on all of them—how does it jointly aggregate information from these three heads to produce a final answer? Second, we observe that heads implement a “self-correction” mechanism indicated by the negative correlation between examples signals, but the exact mechanism of this self-correction remains to be discovered.
Our findings demonstrate how a focused study can reveal sophisticated latent structures and delicate computational strategies within a LLM. We hope this motivates more work to dive deep into the LLM inner structures and obtain more understanding of LLM latent spaces and computations.
Acknowledgments
We thank Lisa Dunlap, Alex Pan, Tony Lian, Anya Ji, Jiaxin Ge, Junyi Zhang, Hanlin Zhu and Shu Liu for their helpful feedback and suggestions. KY is supported by the Vitalik Buterin Ph.D. Fellowship in AI Existential Safety. MJ is supported by the European Union (ERC-2022-SYG-OCEAN-101071601). LC is supported by a Miller Research Fellowship. JS is supported by Open Philanthropy, Simons Foundation, and NSF.
Appendix A Additional table in §3.2
Layer Accuracy
No intervention 0.87
$[0,31]$ 0.85
$[0,15]$ 0.83
$[13,15]$ 0.83
$\{13,15\}$ 0.83
$\{14,15\}$ 0.69
$\{13,14\}$ 0.25
$\{15\}$ 0.71
$\{13\}$ 0.27
$\{14\}$ 0.03
$[0,31]\setminus\{13,15\}$ 0.05
Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers. We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.8673$, while other combinations lead to negligible drops in accuracy, especially when mean-ablating the layer $13$ and $15$ (in blue).
Appendix B Additional figures in §4.3
Figure 10: Head (15,2) projected onto parity direction
Figure 11: Head (15,2) projected out of parity direction
Figure 12: Head (15,2) projected onto mag subspace
Figure 13: Head (15,2) projected out of mag subspace

### v2

@@
 Our results demonstrate how tracking low-dimensional subspaces across a forward pass can provide insight into fine-grained computational structures.
-footnotetext: Equal advising.
+footnotetext: Equal advising.\$*\$\$*\$footnotetext: Correspondence to: Xinyan Hu <xinyanhu@berkeley.edu>.
 1 Introduction
 Large language models (LLMs) have demonstrated remarkable in-context learning (ICL) capabilities, where they can adapt to various new tasks through few-shot demonstrations in the prompt.
-Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; von Oswald et al., 2023; Bai et al., 2023; Zhang et al., 2023; Nichani et al., 2024) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024).
+Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; Akyürek et al., 2023; von Oswald et al., 2023) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024).
 However, little is known about the fine-grained computational structure of in-context learning for large models.
@@
 We first discover that only a small number of model components are responsible for add-k: out of 1024 attention heads, only three do most of the work (§3).
-We find the task information for each head (i.e., the number $k$) is encoded in a six-dimensional subspace; zooming into one of the heads, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4).
+We find the task information (i.e., the number $k$) for each head is encoded in a six-dimensional subspace; zooming into each head, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4).
 We also examine how this head aggregates information from individual in-context examples (§5).
@@
 Since these three heads are responsible for computing add-k, they must extract signal about $k$ from individual examples and then represent $k$ in the residual stream.
-To understand how $k$ is represented, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1).
-Furthermore, zooming into the six-dimensional subspace of one head, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions (with frequencies $2,5,10$), and a two-dimensional subspace tracking the coarse-grained magnitude (the ten digit) (§4.2-4.3).
+To understand how $k$ is represented, zooming into each head, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1).
+Furthermore, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions, and a two-dimensional subspace tracking the coarse-grained magnitude (§4.2-4.3).
 This offers insight on how LLMs internally represent addition, aligning with recent findings showing that (non-ICL) addition is encoded via trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025).
 Finally, we examine how this head aggregates information from individual in-context examples.
-We find that the information is mostly extracted from the $y_{i}$ token positions and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2).
+We find that the information is mostly extracted from the $y_{i}$ token positions, and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2).
 Moreover, we identify a self-correction mechanism: if the signal extracted from one token has an error, later tokens are likely to write in the opposite direction of the error (§5.3).
@@
 2.1 Model and Task
-We begin by specifying the model and task studied in this paper.
+We begin by specifying the model and the task studied in this paper.
 Model.
@@
 For a constant $k$, the add-k task is to add $k$ to a given input integer $x$ to predict $y=x+k$.
-In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$" (where each $y_{i}=x_{i}+k$).
-The examples are concatenated in the prompt using the separator “##\##", followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt.
+In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$” (where each $y_{i}=x_{i}+k$).
+The examples are concatenated in the prompt using the separator “##\##”, followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt.
 A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e.
@@
   3. 3.
-    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$" for some $x_{q}$.
+    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$” for some $x_{q}$.
 2.2 Activation Patching and Function Vectors
@@
 Focusing on two layers via layer-wise ablation.
-To efficiently narrow down the important heads, we first perform ablations at the level of layers.
+To efficiently narrow down the important heads, we first perform mean-ablations at the level of layers.
 From Figure 3(a), we observe that the significant heads $\mathcal{H}_{\text{sig}}$ are concentrated primarily in the middle and late layers.
 We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query.
-After trying different sets of layers, we found that ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while ablating any other combination of layers causes negligible drops in accuracy (Table 3).
+After trying different sets of layers, we found that mean-ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while mean-ablating any other combination of layers causes negligible drops in accuracy (Appendix A, Table 3).
 After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
@@
 No intervention  N/A   0.87
-$(15,2)\text{:=Head1}$ $6$   0.85
-$(15,1)\text{:=Head2}$ $5$   0.83
-$(13,6)\text{:=Head3}$ $5$   0.66
+[…]
+[…]
+[…]
 Any other head  Optimal   0.19
@@
 For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL examples.
-Our next goal is to understand how task information is represented in the activation space.
+Our next goal is thus to understand how task information is represented in the activation space.
 Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
@@
 Thus, we confine our study to this concise six-dimensional subspace for each head.
-The rest of the section focuses on head 1 for simplicity.
+The rest of the section focuses on head 1 and we defer analogous results for head 2 and 3 to appendix B.
 Figure 4: The explained variance ratio vs. number of PCs for each head. The first six PCs make up most of the explained variance ($97\%$) for each head.
@@
 Leveraging the feature directions identified previously, we decompose the head activation subspace into lower-dimensional components that separately encode different subsignals relevant to the task—in this case, the units digit and tens digit.
-By construction, each of the feature directions (viewed as linear projections of $h_{k}$) are periodic functions of $k$.
+By construction, the coordinate function of each feature directions (viewed as linear projections of $h_{k}$) is a periodic function of $k$.
 Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo $T$”.
-Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
+Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
 We verify these hypotheses through causal intervention.
 In Figure 7, we validate hypothesis (ii): we establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal.
-We include analogous experimental results for the two other hypotheses in Appendix B.
+We include analogous experimental results for the two other hypotheses of head 1 in Appendix B.2.
 Figure 7: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto the “unit subspace” (left) and out of the “unit subspace” (right).
-Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully accounts for the final-answer errors.
+Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully account for the final-answer errors.
 This confirms that the unit subspace specifically encodes the unit-digit signal.
@@
 Now, we analyze how the model extracts the task information from the ICL examples.
-In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior.
+In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting task information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior.
 For simplicity, we focus on head 1 in this section.
@@
 We next examine what specific information is extracted from each of these tokens.
-To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},h_{k}\rangle$ between the extracted information and the head vector for each task $k$.
+To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},\tilde{h}_{k}\rangle$ between the extracted information and the head vector (projected onto the subspace and normalized to have unit norm) for each task $k$.
 In Figure 9, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$.
 We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
-Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector $h_{k}$ (projected onto the subspace and normalized to have unit norm).
+Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector (projected onto the subspace and normalized to have unit norm) $\tilde{h}_{k}$.
 For each example $x_{i}\rightarrow y_{i}$, the inner product peaks for $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
@@
 This suggests a self-correction mechanism: intuitively, when the head extracts a noisy signal from one example, signals from subsequent examples are more likely to correct the error, thereby stabilizing the final representation.
-Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlation
+Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlations
 Average $-2.01$ $0.27$
@@
 6 Related Work
-Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Elhage et al., 2021; Wu et al., 2023; Wang et al., 2022; Feng & Steinhardt, 2023; Gould et al., 2023).
+Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Olah et al., 2020; Wang et al., 2022; Wu et al., 2023; Cunningham et al., 2023; Feng & Steinhardt, 2023; Gould et al., 2023; Park et al., 2024).
+Below we discuss three particular threads that are most relevant to this paper.
 Interpreting arithmetic tasks.
-A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Nikankin et al., 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025).
+A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Hanna et al., 2023; Nikankin et al., 2024; Maltoni & Ferrara, 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025).
 Zhou et al.
 (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers.
-Different from them, our work is the first to analyze addition for in-context learning in a large pretrained model, and we find similar representation structures to Zhou et al.
+Different from them, our work is the first to analyze addition for in-context learning in LLM, and interestingly we find similar representation structures to Zhou et al.
 (2024) and Kantamneni & Tegmark (2025).
 Interpreting in-context learning.
-Researchers have constructed detailed models of in-context learning (ICL) for small transformer models (Garg et al., 2022; von Oswald et al., 2023; Bai et al., 2023; Zhang et al., 2023; Nichani et al., 2024), as well as coarser-grained treatments for large pretrained models, attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023).
+Researchers have constructed detailed models of in-context learning (ICL) for small transformer models in standard supervised learning problems such as linear regression (Garg et al., 2022; Akyürek et al., 2023; Zhang et al., 2023; Li et al., 2023; Wu et al., 2024), as well as more complex settings (von Oswald et al., 2023; Bai et al., 2023; Bietti et al., 2023; Reddy, 2023; Guo et al., 2023; Nichani et al., 2024).
+For large pretrained models, there exist coarser-grained treatments attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023).
 Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads.
-Motivated by this, we perform a detailed analysis on the mechanisms behind a family of ICL addition tasks through a novel optimization method inspired by function vectors, which achieves better performance than the method in Todd et al.
+Motivated by this, we study function vector heads in details for a family of ICL tasks, introducing a novel optimization method, which achieves better performance than the method in Todd et al.
 (2024).
-Unlike the tasks studied in Todd et al.
-(2024), our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.
+Another difference from Todd et al.
+(2024) is that our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.
 Causal analysis.
-Our methodological approach draws on recent developments in the causal analysis of neural networks that aims to understand the causal influence of model components to its behavior by performing targeted interventions on the internal states of models (Geiger et al., 2021; Vig et al., 2020).
-In particular, our methods draw inspiration from the causal mediation analysis used in Todd et al.
+There has been a line of research that proposes methods to understand the causal influence of model components on model behavior, such as by probing (Conneau et al., 2018; Hewitt & Manning, 2019; Clark et al., 2019).
+Our methodological approach follows recent developments in revealing causal effects of model components by interventions on internal states of models (Vig et al., 2020; Geiger et al., 2021).
+In particular, we draw inspiration from causal mediation analysis used in Todd et al.
 (2024), activation patching (Meng et al., 2022), and causal scrubbing (Chan et al., 2022).
@@
 Appendix A Additional table in §3.2
+In §3.2, we did systematic ablation studies to narrow down the tasks to three main heads from $33$ significant heads.
+The first step of the ablation studies is layer-wise ablation, where we mean-ablate the significant heads in a subset of layers.
+We include the experimental results here, which narrow down to layer 13 and 15.
 Layer Accuracy
@@
 Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers.
-We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.8673$, while other combinations lead to negligible drops in accuracy, especially when mean-ablating the layer $13$ and $15$ (in blue).
-Appendix B Additional figures in §4.3
-Figure 10: Head (15,2) projected onto parity direction
-Figure 11: Head (15,2) projected out of parity direction
-Figure 12: Head (15,2) projected onto mag subspace
-Figure 13: Head (15,2) projected out of mag subspace
+We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.87$, while other combinations lead to substantial drops in accuracy, especially when mean-ablating layers $13$ and $15$ (in blue).
+Appendix B Additional figures in §4
+In §4.2 and §4.3, we focused on head 1 for simplicity of exposition. Here we include the analogous experimental results for head 2 and 3.
+B.1 Additional figures for §4.2
+In §4.2, we found six trigonometric functions that can be linearly fitted by the coordinate functions of head 1 (Figure 6). Here we show analogous experimental results for head 2 and 3.
+Figure 10: The coordinate functions (inner product with feature directions) of head 2’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
+Figure 11: The coordinate functions (inner product with feature directions) of head 3’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
+B.2 Additional figures for §4.3
+In §4.3, we showed experimental results validating the hypothesis (ii) for head 1 (Figure 7).
+Here we show analogous experimental results validating the hypotheses (i) and (iii) for head 1, as well as all three hypotheses for heads 2 and 3.
+Recall that the three hypotheses are as follows:
+  1. (i)
+    the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task;
+  2. (ii)
+    the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$;
+  3. (iii)
+    the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
+We first show the experimental results validating hypotheses (i) and (iii) for head 1.
+Projecting out of the parity direction doesn’t lead to high errors for the parity and the final answer across all tasks, which might be because parity is relatively easy to obtain (e.g., random choice leads to 0.5 accuracy).
+(a) Head 1 projected onto parity direction
+(b) Head 1 projected out of parity direction
+Figure 12: Validation of hypothesis (i) for head 1.
+(a) Head 1 projected onto magnitude subspace
+(b) Head 1 projected out of magnitude subspace
+Figure 13: Validation of hypothesis (iii) for head 1.
+We then show analogous results for head 2.
+(a) Head 2 projected onto parity direction
+(b) Head 2 projected out of parity direction
+Figure 14: Validation of hypothesis (i) for head 2.
+(a) Head 2 projected onto unit subspace
+(b) Head 2 projected out of unit subspace
+Figure 15: Validation of hypothesis (ii) for head 2.
+(a) Head 2 projected onto magnitude subspace
+(b) Head 2 projected out of magnitude subspace
+Figure 16: Validation of hypothesis (iii) for head 2.
+Finally, we show analogous results for head 3.
+The evidence for head 3 is slightly weaker than the first two heads, which aligns with the fact that the intervened accuracy of head 3 is slightly weaker than the ones of the first two heads (Tabel 1).
+(a) Head 3 projected onto parity direction
+(b) Head 3 projected out of parity direction
+Figure 17: Validation of hypothesis (i) for head 3.
+(a) Head 3 projected onto unit subspace
+(b) Head 3 projected out of unit subspace
+Figure 18: Validation of hypothesis (ii) for head 3.
+(a) Head 3 projected onto magnitude subspace
+(b) Head 3 projected out of magnitude subspace
+Figure 19: Validation of hypothesis (iii) for head 3.

### v3

@@
 Abstract
-To perform in-context learning, language models must extract signals from individual few-shot examples, aggregate these into a learned prediction rule, and then apply this rule to new examples.
+To perform few-shot learning, language models extract signals from a few input-label pairs, aggregate these into a learned prediction rule, and apply this rule to new inputs.
 How is this implemented in the forward pass of modern transformer models?
-To study this, we consider a structured family of few-shot learning tasks for which the true prediction rule is to add an integer $k$ to the input.
-We find that Llama-3-8B attains high accuracy on this task for a range of $k$, and localize its few-shot ability to just three attention heads via a novel optimization approach.
-We further show the extracted signals lie in a six-dimensional subspace, where four of the dimensions track the unit digit and the other two dimensions track overall magnitude.
-We finally examine how these heads extract information from individual few-shot examples, identifying a self-correction mechanism in which mistakes from earlier examples are suppressed by later examples.
-Our results demonstrate how tracking low-dimensional subspaces across a forward pass can provide insight into fine-grained computational structures.
+To explore this question, we study a structured family of few-shot learning tasks for which the true prediction rule is to add an integer $k$ to the input.
+We introduce a novel optimization method that localizes the model’s few-shot ability to only a few attention heads.
+We then perform an in-depth analysis of individual heads, via dimensionality reduction and decomposition.
+As an example, on Llama-3-8B-instruct, we reduce its mechanism on our tasks to just three attention heads with six-dimensional subspaces, where four dimensions track the unit digit with trigonometric functions at periods $2$, 55, and $10$, and two dimensions track magnitude with low-frequency components.
+To deepen our understanding of the mechanism, we also derive a mathematical identity relating “aggregation” and “extraction” subspaces for attention heads, allowing us to track the flow of information from individual examples to a final aggregated concept.
+Using this, we identify a self-correction mechanism where mistakes learned from earlier demonstrations are suppressed by later demonstrations.
+Our results demonstrate how tracking low-dimensional subspaces of localized heads across a forward pass can provide insight into fine-grained computational structures in language models.
 footnotetext: Equal advising.\$*\$\$*\$footnotetext: Correspondence to: Xinyan Hu <xinyanhu@berkeley.edu>.
 1 Introduction
-Large language models (LLMs) have demonstrated remarkable in-context learning (ICL) capabilities, where they can adapt to various new tasks through few-shot demonstrations in the prompt.
-Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; Akyürek et al., 2023; von Oswald et al., 2023) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024).
-However, little is known about the fine-grained computational structure of in-context learning for large models.
-To address this, we perform a detailed analysis of ICL for a structured family of tasks: learning to add a constant number (“add-k”).
-For these tasks, models are given an input sequence “$x_{1}\rightarrow y_{1}\#...\#x_{n}\rightarrow y_{n}\#x_{q}\rightarrow$” where $y_{i}=x_{i}+k$ for all $i$.
-This task is mathematically well-defined and tractable for an in-depth mechanistic analysis, while exhibiting the core challenge of extracting and applying a pattern from ICL examples.
-Moreover, modern language models are empirically capable of solving it (they correctly predict $y_{q}=x_{q}+k$).
-Figure 1: In this paper, we study the activation space of attention heads in Llama-3-8B during in-context learning (ICL) for addition tasks.
-Our main findings: (1) out of $1024$ attention heads in Llama-3-8B, only three are important for these tasks (§4); (2) each head encodes the task information in a six-dimensional subspace (§4.1); (3) a six-dimensional subspace further decomposes into a four-dimensional subspace capturing the unit digit of the task information and a two-dimensional subspace capturing its ten digit (§4.3); and (4) task information is mostly extracted from a subset of tokens (§5).
-The figure illustrates these findings using an example add-k prompt.
-To understand the mechanism behind in-context add-k, we study it in Llama-3-8B.
-We first discover that only a small number of model components are responsible for add-k: out of 1024 attention heads, only three do most of the work (§3).
-We find the task information (i.e., the number $k$) for each head is encoded in a six-dimensional subspace; zooming into each head, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4).
-We also examine how this head aggregates information from individual in-context examples (§5).
-To discover the attention heads that are responsible for add-k, we introduce a novel approach inspired by previous work on function vectors (Todd et al., 2024), which are vectors that imitate ICL when they are patched into the residual stream (§2.2).
-We formulate an optimization method that identifies $33$ important heads whose outputs yield good ICL performance (§3.1).
-Through further ablation, we find that all but 3 of these 33 heads can be replaced with their mean value without significantly decreasing the performance (§3.2).
-Since these three heads are responsible for computing add-k, they must extract signal about $k$ from individual examples and then represent $k$ in the residual stream.
-To understand how $k$ is represented, zooming into each head, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1).
-Furthermore, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions, and a two-dimensional subspace tracking the coarse-grained magnitude (§4.2-4.3).
-This offers insight on how LLMs internally represent addition, aligning with recent findings showing that (non-ICL) addition is encoded via trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025).
-Finally, we examine how this head aggregates information from individual in-context examples.
-We find that the information is mostly extracted from the $y_{i}$ token positions, and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2).
-Moreover, we identify a self-correction mechanism: if the signal extracted from one token has an error, later tokens are likely to write in the opposite direction of the error (§5.3).
-This means that ICL goes beyond simple averaging of inputs and has stateful dynamics.
-Our findings suggest that ICL tasks can be implemented through specialized activation subspaces from a small number of attention heads that extract, represent and aggregate task-related information in interpretable ways.
-We introduce a novel optimization method to identify important attention heads and demonstrate how a focused analysis of a simple ICL task can reveal delicate latent space structures and sophisticated computational strategies in a large language model.
+Large language models (LLMs) exhibit in-context learning (ICL) abilities; for instance, they can few-shot learn new tasks from a small number of demonstrations in the prompt.
+To understand this ability, past works have constructed detailed models of ICL for small synthetic language models (Garg et al., 2022; Akyürek et al., 2023; von Oswald et al., 2023) as well as coarser-grained analyses of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024).
+However, little is known about the fine-grained computational structure of ICL for large models.
+ICL extracts task information from demonstrations and applies the aggregated information to input queries.
+Previous work (Todd et al., 2024) constructed a vector (i.e., the “function” vector) from demonstrations that encode the task information—for instance, adding it to the residual stream on zero-shot inputs recovers ICL behavior.
+However, two questions remain elusive: (1) How precisely do function vectors encode task information?
+(2) How do models aggregate information from ICL examples to form such function vectors?
+To address these questions, we perform a case study for few-shot learning of arithmetic (i.e., learning to add a constant $k$ to the input).
+This family of tasks has the advantage of providing a large number of tasks (different integers $k$) that all share the same input domain, which is important to rule out domain-based shortcuts when analyzing ICL mechanisms (§2.1).
+Arithmetic also has the advantage of being well-studied in other (non-in-context) settings (Zhou et al., 2024; Kantamneni & Tegmark, 2025), allowing us to situate our results with other known mechanisms.
+Finally, most LLMs perform this task reliably (e.g.
+$87\%$ accuracy for Llama-3 and $90\%$ for Qwen-2.5).
+To study this setting, we first introduce a novel optimization method for localizing few-shot ability to a small number of attention heads.
+Our approach is inspired by previous work on function vectors (Todd et al., 2024), which mimic ICL when they are patched into the residual stream (§2.2).
+While Todd et al.
+(2024) selected heads based on their individual effect on ICL performance, we search for coefficients within a continuous range $[0,1]$ that produce a sparse, weighted combination of heads’ outputs, maximizing ICL performance (§3.1).
+Our method successfully finds a small number of heads that recover most of the ICL performance (§3.2); for example, Llama-3-8B-instruct achieves 79% accuracy using function vectors from just three heads (90% of the original ICL accuracy).
+We next perform a detailed analysis of these heads, through dimensionality reduction and decomposition.
+We find that the task information from each head is encoded in a low-dimensional subspace typically consisting of trigonometric functions.
+For example, in each of the three important heads in Llama-3-8B-instruct, we get a 6-dimensional subspace, where 4 dimensions are periodic functions with periods 2, 5, and 10 (tracking the units digit) and 2 dimensions are low-frequency functions (tracking the tens digit) (§4).
+This interestingly aligns with recent findings in non-in-context settings where addition is also encoded by trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025), suggesting a deeper relation between non-ICL and ICL machinery.
+Figure 1: Key findings of our methods in the specific case of Llama-3-8B-instruct (illustrated using an example add-k prompt): (1) out of 1024 attention heads, only three are important for add-k ICL (§4); (2) each head encodes the task information $k$ in a six-dimensional subspace (§4.1); (3) the six-dimensional subspace further factors into four dimensions capturing the unit digit of $k$ (encoding periodic functions at periods 2,5, and 10) and two dimensions capturing its tens digit (encoding higher frequency functions) (§4.3); and (4) task information flows from output tokens to final token via a self-correction mechanism (§5).
+Finally, we further study the flow of information across tokens, by deriving a general mathematical relation between “extractor” and “aggregator” subspaces, building on the second-order logit lens (Gandelsman et al., 2025).
+This lets us study the task-related information that heads extract from each token.
+For example, in Llama-3-8B-instruct we find that task information is mostly extracted from label tokens (§5.2).
+Moreover, we find a self-correction mechanism: if the signal extracted from one token has an error, signals from later tokens often write in the opposite direction of the error (§5.3).
+This suggests that ICL goes beyond simple averaging of inputs and has stateful dynamics.
+In summary, we found that task-specific information in ICL can be localized to a small number of attention heads and low-dimensional subspaces with structured patterns.
+We also found that models employ a self-correction mechanism when the task information flows from the “extractor” subspaces to the “aggregator” subspaces.
+Our findings show how even in large models, ICL mechanisms can be localized to specialized activation subspaces from a small number of heads that extract, represent, and aggregate information in interpretable ways.
+Our key contributions are thus as follows: (1) we introduce a novel optimization method to identify relevant attention heads for ICL; (2) we derive a precise mathematical relation between signals from earlier tokens and the output token; and (3) we perform a focused analysis of ICL addition and are able to reverse-engineer rich latent structures and sophisticated computational strategies in LLMs.
 2 Preliminaries
@@
 Model.
-We study Llama3-8B-Instruct, which has 32 layers and 32 attention heads per layer, and a residual dimension of 4096.
+We focus on Llama-3-8B-instruct in the main body, which has 32 layers and 32 attention heads per layer and a residual dimension of 4096.
 We denote each head as a tuple (layer index, head index), where both indices range from 0 to 31.
+To make our analysis broader in model size, training type, and model family, we also experiment on Llama-3.2-3B-instruct, Llama-3.2-3B, and Qwen-2.5-7B and report the results in appendix.
+The results are consistent across Llama family and directionally similar for the Qwen model.
 Task.
+We study a structured family of ICL tasks that we call add-$k$.
 For a constant $k$, the add-$k$ task is to add $k$ to a given input integer $x$ to predict $y=x+k$.
-In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$” (where each $y_{i}=x_{i}+k$).
-The examples are concatenated in the prompt using the separator “##\##”, followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt.
+In an $n$-shot ICL prompt, the model is given $n$ demonstrations of the form $x_{i}\rightarrow y_{i}$” with $y_{i}=x_{i}+k$, concatenated using the separator “#\#”, followed by a query $x_{q}\rightarrow$” (see Figure 1 for an example).
 A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e.
-$y_{i}-x_{i}$ in the prompts) but not the input domain, in contrast to most previous studies (Todd et al., 2024; Hendel et al., 2023).
-We construct data for the task by varying $x\in[1,100]$ and $k\in[1,30]$ (thus $y\in[2,130]$) since Llama3-8B is empirically capable of solving the addition task in this range.
+$y_{i}-x_{i}$) but not in the input domain, enabling us to isolate task information from input content so as to dissect the ICL mechanism at a finer granularity.
+[…]
+We construct data for the task by varying $x\in[1,100]$ and $k\in[1,30]$ (thus $y\in[2,130]$) since Llama-3 models are empirically capable of solving the addition task in this range.
 We consider the following three types of prompts:
   1. 1.
-    five-shot ICL prompt, where all 5 demonstration examples satisfy $y_{i}=x_{i}+k$ for a fixed $k\in[1,30]$. We also call this add-k ICL.
+    five-shot ICL prompt, where all five demonstrations satisfy $y_{i}=x_{i}+k$ for a fixed $k\in[1,30]$. We also call this add-k ICL.
   2. 2.
-    mixed-$k$ ICL prompt, where the five demonstration examples are $y_{i}=x_{i}+k_{i}$ for five possibly different $k_{i}$ values.
+    mixed-$k$ ICL prompt, where the demonstrations are $y_{i}=x_{i}+k_{i}$ for possibly different $k_{i}$ values.
   3. 3.
-    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$” for some $x_{q}$.
+    zero-shot prompt, where there are no demonstrations: the prompt is $x_{q}\rightarrow$” for some $x_{q}$.
+We mostly study five-shot ICL prompts as examples of our ICL tasks, and use them to generate function vectors (§2.2).
+We also examine the information extracted from demonstrations in mixed-$k$ ICL prompts, which is a varied version of five-shot ICL prompts with mistaken demonstrations, in §5.2.
+We use zero-shot prompts to evaluate the effectiveness of heads and function vectors (§2.2).
 2.2 Activation Patching and Function Vectors
-Next, we briefly introduce activation patching, a common interpretability technique that will be used throughout the paper, and the function vector, a construction that we will use to identify important heads for our ICL tasks.
+Next, we briefly review activation patching, a common interpretability technique that will be used throughout the paper, and the function vector, a construction that we will use to identify important heads for our ICL tasks.
 Activation patching.
@@
 Specifically, if $z_{l}$ is the original value of the residual stream at layer $l$ at the final token position $\to$”, we patch in the replacement $z_{l}+v$, where $v$ is constructed from activations on a different input prompt.
-We choose the layer $l$ at one third of the network’s depth and construct $v$ from “function vector” heads, following Todd et al.
+We choose the layer $l$ at one-third of the network’s depth and construct $v$ from “function vector” heads, following Todd et al.
 (2024), as described next.
@@
 We define the intervention accuracy of $v_{k}$ as the average accuracy obtained on zero-shot prompts when $v_{k}$ is added to the residual stream across all tasks $k$.
-This metric captures how effectively $v_{k}$ encodes task-specific information about $k$.
+Due to the independency of $k$ from input queries, this metric captures how effectively $v_{k}$ encodes task-specific information about $k$.
 For comparison, we define clean accuracy as the accuracy on five-shot prompts without any intervention, also averaged across all tasks $k$.
@@
 [^1] Similar to Todd et al.
-(2024), we approximate this average using $100$ randomly generated five-shot add-k prompts, where each prompt contains random demonstration examples $x_{i}$ and the final query input $x_{q}$ is chosen exactly once from each integer in $[1,100]$.
+(2024), we approximate this average using $100$ randomly generated five-shot add-k prompts, where each prompt contains random demonstration inputs $x_{i}$ and the final query input $x_{q}$ is chosen exactly once from each integer in $[1,100]$.
 Todd et al.
@@
 Building on this framework, we consider multiple ways to construct $v_{k}$: (1) the task-specific mean $h_{k}$ over the add-k task (as described above); (2) the overall mean $\bar{h}$ (i.e., average across all $k$), which removes task-specific information about $k$; or (3) the specific value $h(p)$ on an individual prompt $p$.
-Throughout the paper, we call $\bar{h}$ the mean-ablation of head $h$ and call $h_{k}$ the head vector of $h$.
+Throughout the paper, we call $\bar{h}$ the mean-ablation of head $h$ and call $h_{k}$ the head vector of $h$ with respect to $k$.
 Beyond the three choices above, we sometimes project a head’s output onto a lower-dimensional subspace or scale it by a coefficient.
@@
 Figure 2: The chain of localization in §3 and §4.
-We first identify $33$ significant attention heads (out of 1024) via a global optimization method (§3.1), then narrow down to 3 main heads while mean-ablating the remaining 30 heads (§3.2).
-We next study the structure of the representations of each main head by localizing it to a six-dimensional subspace (§4.1), and decompose it into a four-dimensional subspace encoding the unit digit and a two-dimensional subspace encoding the tens digit (§4.3).
+We first identify $33$ significant attention heads (out of 1024) via a global optimization method (§3.1), then narrow down to 3 main heads while mean-ablating the remaining 30 (§3.2).
+We next study the structure of the representation of each main head by localizing it to a six-dimensional subspace (§4.1), and decompose it into a four-dimensional subspace encoding the unit digit and a two-dimensional subspace encoding the tens digit (§4.3).
 3 Identifying Three Aggregator Heads
@@
 In this section, we find that three heads do most of the work for add-k.
-We find these heads by first solving an optimization problem to identify $33$ significant heads out of 10241024102410243.1), and then further narrowing down to three main heads via systematic ablations (§3.2).
+We find these heads by first solving an optimization problem to identify $33$ significant heads out of 102410243.1), and then further narrowing down to three main heads via systematic ablations (§3.2).
 This process is illustrated in the first row of Figure 2.
@@
 (Recall that $h_{k}$ is the average output of head $h$ averaged across a large dataset of five-shot add-k prompts.) We will search for a sparse coefficient vector $c\in[0,1]^{1024}$ such that adding $v_{k}(c)$ to the residual stream of a zero-shot prompt achieves high accuracy on the add-k task.
-Let $\ell(x_{q},y_{q};v)$ be the cross-entropy loss between $y_{q}$ and the model output when intervening the vector $v$ onto the input “$x_{q}\to$”(i.e.
-replacing $z_{l}$ with $z_{l}+v$ in the forward pass on “$x_{q}\to$”, where $z_{l}$ is the layer-$l$ residual stream at the last token).
+Let $\ell(x_{q},y_{q};v)$ be the cross-entropy loss between $y_{q}$ and the model output when intervening the vector $v$ onto the input $x_{q}\to$”(i.e.
+replacing $z_{l}$ with $z_{l}+v$ in the forward pass on $x_{q}\to$”, where $z_{l}$ is the layer-$l$ residual stream at the last token).
 We optimize $c$ with respect to the loss
-$\mathcal{L}(c)=\mathbb{E}_{k\in[30]}\mathbb{E}_{x_{q}\in[100]}[\ell(x_{q},x_{q}+k;v_{k}(c))]+\lambda\|c\|_{1},$ (1)
+$\mathcal{L}(c)=\mathbb{E}_{k,x_{q}}[\ell(x_{q},x_{q}+k;v_{k}(c))]+\lambda\|c\|_{1},$ (1)
 where the regularization term with weight $\lambda$ promotes sparsity.
-(a) Coefficients of heads.
-(b) Average indirect effects of heads (Todd et al., 2024).
-Figure 3: Comparison of significant heads identified by our optimized coefficients (left) and by average indirect effects (AIE) from the previous method (Todd et al., 2024) (right).
-Colors indicate the magnitude of each head’s importance (coefficients or AIE).
-The top $33$ heads identified by both methods are highlighted with frames ($13$ heads common across both methods in red and other $20$ heads in blue).
-Our identified heads yield an intervention accuracy of $0.85$, compared to the previous method’s accuracy of $0.31$.
-Both methods select heads from similar layers, but our optimization approach is significantly more effective.
 Training details.
-We randomly select $25$ add-k tasks from our total $30$ tasks for training and in-distribution testing, and use the remaining five tasks only for out-of-distribution testing.
-For each task add-k, we generate $100$ zero-shot prompts “$x_{q}\rightarrow$”, where $x_{q}$ ranges over all integers from $[1,100]$ and the target output is $x_{q}+k$, yielding one data point for each (prompt, task) pair.
-We randomly split the data points of the $25$ tasks into training, validation, and test sets in proportions of $0.7$, $0.15$, and $0.15$, respectively.
+We randomly select $25$ add-k tasks of the total $30$ tasks for training and in-distribution testing, and use the remaining five tasks only for out-of-distribution testing.
+For each task add-k, we generate $100$ zero-shot prompts $x_{q}\rightarrow$”, where $x_{q}$ ranges over all integers from $[1,100]$ and the target output is $x_{q}+k$, yielding one data point for each (prompt, task) pair.
+We randomly split the data points of the $25$ tasks into training, validation, and test sets in proportions $0.7$, $0.15$, and $0.15$, respectively.
 We use AdamW with learning rate $0.01$ and batch size $128$.
@@
 Results.
-We get coefficients that achieves high intervention accuracy.
-In particular, at the final epoch, the average accuracies of intervening on test data for the $25$ tasks and the five tasks ($0.83$ and $0.87$) are both close to the average accuracies of doing no intervention ($0.89$ and $0.92$).
+We get coefficients that achieve high intervention accuracy.
+In particular, on the $25$ in-distribution tasks and five out-of-distribution tasks, intervention accuracies at the final epoch are $0.83$ and $0.87$, close to the clean accuracies of $0.89$ and $0.92$, respectively.
 To identify the important heads for the tasks, we plot the values of the coefficients in the final epoch for each layer and head index (Figure 3(a)).
@@
 Using this construction, we achieve an intervention accuracy of $0.85$, close to the clean accuracy of $0.87$, indicating that our $33$ heads captures most of the necessary information for the add-k task.
-In contrast, selecting the top 33 heads according to AIE yields a much lower intervention accuracy of $0.31$.
-[^2] Todd et al.
+In contrast, selecting the top 33 heads according to AIE yields a much lower intervention accuracy of 0.310.31.222Todd et al.
 select the top 10 heads in their work, but this yields an even lower accuracy of $0.05$ in our setting.
-In Figure 3, we visualize the coefficients and AIE values of heads from both methods and highlight the $13$ heads that are common to both approaches.
+We visualize the coefficients and AIE values of heads from both methods in Figure 3.
 3.2 Further Refinement via Ablations
@@
 Intuitively, while we require the overall signal transmitted by these heads, we do not need any information about the specific value of $k$.
-To test this hypothesis, we perform mean-ablations: replacing each task-specific signal $h_{k}$ with the overall mean $\overline{h}$ across all values of $k$2.2).
+To test this hypothesis, we perform mean-ablations: replacing each task-specific signal $h_{k}$ with the overall mean $\overline{h}$ across all values of kk2.2).
 Specifically, we conduct mean-ablations over subsets of the 33 significant heads and measure the resulting intervention accuracy of the corresponding function vectors.
 Formally, when ablating a subset $\mathcal{H}_{0}$, the resulting function vector is given by
-$v=\sum_{h\in\mathcal{H}_{0}}\overline{h}+\sum_{h\in\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}}h_{k}.$ (2)
+$v_{k}=\sum_{h\in\mathcal{H}_{0}}\overline{h}+\sum_{h\in\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}}h_{k}.$ (2)
 Here, heads in $\mathcal{H}_{0}$ contribute only their overall mean signals, while heads in $\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}$ retain their task-specific information.
@@
 We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query.
-After trying different sets of layers, we found that mean-ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while mean-ablating any other combination of layers causes negligible drops in accuracy (Appendix A, Table 3).
+After trying different sets of layers, we found that mean-ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while mean-ablating any other combination of layers causes negligible drops in accuracy (Appendix B.1, Table 3).
 After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
@@
 To amplify each head’s effect, we scale its output by a coefficient (e.g., 5).
-We find that three heads—head 1 = $(15,2)$, head 2 = $(15,1)$, and head 3 = $(13,6)$—achieve intervention accuracies close to the clean accuracy when appropriately scaled, while all other heads show much lower accuracies regardless of scaling (Table 1).
+We find that three heads—head 1 = $(15,2)$, head 2 = $(15,1)$, and head 3 = (13,6)(13,6)—achieve intervention accuracies close to the clean accuracy when appropriately scaled, while all other heads show much lower accuracies regardless of scaling (Table 1).
 This suggests that these three heads individually encode the task information much better than any others.
@@
 No intervention  N/A   0.87
-[…]
-[…]
-[…]
-Any other head  Optimal   0.19
+$(15,2)\text{:=Head 1}$ 66   0.85
+$(15,1)\text{:=Head 2}$ 55   0.83
+$(13,6)\text{:=Head 3}$ 55   0.66
+Any other head  Optimala   0.19
+  • a
+    Coefficient chosen to maximize accuracy for each head (scanned over integer values).
 Table 1: Intervened accuracies for scaling up each single head’s output by an optimal coefficient.
@@
 4 Characterizing the Aggregator Subspace
-For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL examples.
+For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL demonstrations.
 Our next goal is thus to understand how task information is represented in the activation space.
 Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
-In this section, we dissect their activation spaces in three stages: (1) Localize a six–dimensional task subspace in each head via principal component analysis (PCA) (§4.1); (2) Rotate this subspace into orthogonal feature directions aligned with sinusoidal patterns across $k$4.2); (3) Decompose the six-dimensional space into a four-dimensional unit-digit subspace and a two-dimensional magnitude subspace that separately encode the units and tens of the answer (§4.3).
+In this section, we dissect their activation spaces in three stages: (1) Localize a six–dimensional task subspace in each head via principal component analysis (PCA) (§4.1); (2) Rotate this subspace into orthogonal feature directions aligned with sinusoidal patterns across kk4.2); (3) Decompose the six-dimensional space into a four-dimensional unit-digit subspace and a two-dimensional magnitude subspace that separately encode the units and tens of the answer (§4.3).
 4.1 Localizing to Six-dimensional Subspace
-To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Figure 4).
+To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Appendix C, Figure 4(a)).
 We then check that the function vectors found earlier remain effective after projecting onto the subspace.
@@
 We find that $\tilde{v}_{k}$ has intervention accuracy $0.76$, which is close to the intervention accuracy of $0.79$ before projection (§3.2).
-Thus, we confine our study to this concise six-dimensional subspace for each head.
-The rest of the section focuses on head 1 and we defer analogous results for head 2 and 3 to appendix B.
-Figure 4: The explained variance ratio vs. number of PCs for each head. The first six PCs make up most of the explained variance ($97\%$) for each head.
-Figure 5: Coordinates of head 1’s vectors (their inner products with a PC) for the first six PCs across different add-$k$ tasks. The coordinates in the first five PCs reveal periodic patterns.
+Thus, we confine our study to the concise six-dimensional subspace for each head.
 4.2 Identifying Feature Directions Encoding Periodic Patterns
-To understand how the six-dimensional subspace of a head represents the task information, we first examine the coordinates of the head vectors (their inner products with the PCs) as a function of $k$.
-This reveals partially periodic patterns in the first five components (Figure 5).
+To understand how the six-dimensional subspace of each head represents task information, we first examine coordinates of head vectors (their inner products with principal components (PC)) as functions of $k$.
+This reveals partially periodic patterns in the first five PCs (Appendix D.1, Figure 5).
 This motivates us to linearly transform the six PCs to find directions that encode pure periodic patterns.
@@
 To find trigonometric functions to fit, we searched over different periods and phases and performed least squares regression.
-We found six trigonometric functions at periods $2$, $5$, $10$, $10$, $25$, and $50$ that could be expressed as functions of the top 6 PCs with low regression error (Figure 6).
+We found six trigonometric functions at periods $2$, 55, $10$, $10$, $25$, and $50$ that could be expressed as functions of the top 6 PCs with low regression error (Appendix D.1, Figure 6).
 We apply the resulting linear transformation to the six PCs to obtain a new set of directions that encode these six pure periodic patterns, which we call feature directions.
-Figure 6: The coordinate functions (inner product with feature directions) of head 1’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
 4.3 Decomposing to Subspaces Encoding Subsignals
 Leveraging the feature directions identified previously, we decompose the head activation subspace into lower-dimensional components that separately encode different subsignals relevant to the task—in this case, the units digit and tens digit.
-By construction, the coordinate function of each feature directions (viewed as linear projections of $h_{k}$) is a periodic function of $k$.
-Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo $T$”.
-Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
+By construction, the coordinate function of each feature direction (viewed as linear projections of $h_{k}$) is a periodic function of $k$.
+Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo TT”.
+Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
 We verify these hypotheses through causal intervention.
-In Figure 7, we validate hypothesis (ii): we establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal.
-We include analogous experimental results for the two other hypotheses of head 1 in Appendix B.2.
-Figure 7: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto the “unit subspace” (left) and out of the “unit subspace” (right).
-Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully account for the final-answer errors.
-This confirms that the unit subspace specifically encodes the unit-digit signal.
-5 Signal Extractors of ICL Examples
-Previously, we localized the model’s behavior to three heads and their corresponding six-dimensional subspaces, then examined how the model represents the task information ($k$ for add-k) inferred from the ICL examples in one subspace.
-Now, we analyze how the model extracts the task information from the ICL examples.
-In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting task information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior.
-For simplicity, we focus on head 1 in this section.
+We establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal.
+We defer experimental results in Appendix E.
+5 Signal Extractors of ICL Demonstrations
+Previously, we localized the behavior to three heads and their six-dimensional subspaces, then examined how the model represents the task information within these subspaces.
+Now, we analyze how the model extracts the task information from the ICL demonstrations.
+In this section, we find that: (1) the signal is primarily gathered from the label tokens in each demonstration; (2) each demonstration $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context demonstrations with conflicting task information; and (3) when all demonstrations $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior.
 5.1 Mathematical Observation: Tracing Subspace back to Previous Tokens
@@
 Thus, the signal extracted from previous tokens is the transformed residual stream at that token.
-Formally, a head $h$’s output at the last token of a prompt $p$ can be written as
-$h(p)=\sum_{t\in p}\alpha_{t}\cdot O_{h}V_{h}\cdot z_{t},$ (3)
+Formally, a head $h$’s output at the last token of a prompt $p$ can be written as $h(p)\;=\;\sum_{t\in p}\alpha_{t}\,O_{h}V_{h}\,z_{t},$
 where $\alpha_{t}$ is the attention score from the last token to each token $t$, satisfying $\sum_{t\in p}\alpha_{t}=1$, $z_{t}$ is the residual stream input to the head $h$ at token $t$, $V_{h}$ is the value matrix, and $O_{h}$ is the output matrix mapping from head-dimensional space to model-dimensional space.
 Let $W_{h}$ denote the projection matrix onto the six-dimensional subspace for head $h$.
-Then the projected signal at the final token, $W_{h}\cdot h(p)$, can be decomposed into contributions from previous tokens as $W_{h}\cdot\alpha_{t}O_{h}V_{h}z_{t}$, each lying in the transformed head subspace defined by the matrix $W_{h}\cdot O_{h}V_{h}$.
-In the following subsection, we analyze the magnitudes and directions of these signals, and study how signals from different examples interact.
-5.2 Signal Extractor for each Example
-Figure 8: For a mixed-$k$ ICL prompt, the strength of the task-signal contribution of each previous token $t$ to the final token, $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$.
-Decomposing it into two parts: (1) the norm of the extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$ (top), and (2) the attention score from the final token $\alpha_{t}$ (bottom), both parts consistently peak at the tokens $t=y_{i}$ (in bold).
-To understand how examples contribute to model generation at the last token, we identify which tokens contribute the most, then examine what information they provide.
+Then the projected signal at the final token, $W_{h}\cdot h(p)$, can be decomposed into contributions from previous tokens as $W_{h}\cdot\alpha_{t}O_{h}V_{h}z_{t}$, each lying in the image of the head subspace under $W_{h}O_{h}V_{h}$.
+In the following subsection, we analyze the magnitudes and directions of these signals, and study how signals from different demonstrations interact.
+5.2 Signal Extractor for each Demonstration
+To understand how demonstrations contribute to model generation at the last token, we identify which tokens contribute the most, then examine what information they provide.
 By the analysis above, the task-signal contribution of each previous token to the final token through the head $h$ is $\alpha_{t}W_{h}O_{h}V_{h}z_{t}$.
 This can be decomposed into two parts: (1) extracted information: $W_{h}O_{h}V_{h}z_{t}$, the residual stream input projected into the relevant subspace; and (2) aggregation weight: $\alpha_{t}$, the attention score of the final token to the previous token.
-We plot the norms of the extracted information and the aggregation weights for a random mixed-$k$ ICL prompt in Figure 8.
+We plot the norms of the extracted information and the aggregation weights for a random mixed-$k$ ICL prompt (with conflicting $k_{i}$ values), in Appendix F.1.
 Both the strength of the extracted information and the aggregation weights peak at $y_{i}$ tokens.
@@
 To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},\tilde{h}_{k}\rangle$ between the extracted information and the head vector (projected onto the subspace and normalized to have unit norm) for each task $k$.
-In Figure 9, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$.
-We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
-Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector (projected onto the subspace and normalized to have unit norm) $\tilde{h}_{k}$.
-For each example $x_{i}\rightarrow y_{i}$, the inner product peaks for $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
-5.3 Signal Correlation among Examples
-Having studied the signal extracted from each individual example in the previous subsection, we next study how signals from different examples interact to execute the ICL task.
-To do so, we compute the correlation between the extracted signals from different demonstration examples: for each $y_{i}$ token, we first compute the inner product between the residual stream input to head 1, $z_{t}$, and the corresponding task vector $h_{k}$, where $k=y_{i}-x_{i}$.
-Then, we compute the correlation of these measures across each pair of five positions over $100$ add-k prompts, yielding $\binom{5}{2}$ correlation values per task.
-To analyze the correlation, we sum the negative correlation values and positive correlation values respectively for each task, and calculate the various statistics (max, average and min) over all tasks (Table 2).
-The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, indicating that the signals from any two demonstration examples are mostly negatively correlated.
-This suggests a self-correction mechanism: intuitively, when the head extracts a noisy signal from one example, signals from subsequent examples are more likely to correct the error, thereby stabilizing the final representation.
-Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlations
-Average $-2.01$ $0.27$
-Min $-1.40$ $0.07$
-Max $-2.34$ $0.54$
-Table 2: The statistics (min, max, and average) of the absolute values of negative correlation sum and positive correlation sum respectively over the $30$ tasks.
-The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, which indicates that the signals from any two demonstration examples are mostly negatively correlated, suggesting a self-correction mechanism.
+In Figure LABEL:fig:15_2_extractor_direction, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$.
+We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding demonstration $x_{i}\rightarrow y_{i}$.
+5.3 Signal Correlation among Demonstrations
+Having studied the signal extracted from each demonstration in the previous subsection, we next study how signals from different demonstrations interact to execute ICL task.
+To do so, we compute the correlation between the extracted signal from different demonstrations: for each $y_{i}$ token, we first compute the inner product between the residual stream input to head 1, $z_{t}$, and the corresponding task vector $h_{k}$, where $k=y_{i}-x_{i}$.
+Then, we compute the correlation of these measures across each pair of five positions over $100$ add-k prompts, yielding $\binom{5}{2}$ correlations per task.
+To analyze the correlation, we sum the negative correlation values and positive correlation values respectively for each task, and calculate the various statistics (max, average and min) over all tasks (Appendix G.1, Table 8).
+The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, indicating that the signals from any two demonstrations are mostly negatively correlated.
+This suggests a self-correction mechanism: intuitively, when the head extracts a noisy signal from one demonstration, signals from subsequent demonstrations are more likely to correct the error, thereby stabilizing the final representation.
 6 Related Work
-Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Olah et al., 2020; Wang et al., 2022; Wu et al., 2023; Cunningham et al., 2023; Feng & Steinhardt, 2023; Gould et al., 2023; Park et al., 2024).
+Our work builds on a growing body of interpretability research that aims to uncover circuits and internal computations of language models.
+Many studies focus on synthetic tasks or models specifically trained on that task (Nanda et al., 2023; Bietti et al., 2023; Reddy, 2023; Singh et al., 2024).
+Going beyond to large pretrained language models, some papers study general LLMs and tasks but provide only coarser-grained analyses (Todd et al., 2024; Olsson et al., 2022; Hendel et al., 2023), while others focus on particular model families and specific task classes to obtain more fine-grained insights (Hanna et al., 2023; Feng & Steinhardt, 2023; Wu et al., 2023; Zhou et al., 2024; Panickssery et al., 2024).
+Our work follows the latter trajectory: we analyze Llama-3 models and the Qwen-2.5 model on a structured set of addition ICL tasks, and we provide a deeper and more detailed account of ICL mechanisms than prior studies of ICL in LLMs (Todd et al., 2024; Olsson et al., 2022; Hendel et al., 2023).
 Below we discuss three particular threads that are most relevant to this paper.
@@
 (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers.
-Different from them, our work is the first to analyze addition for in-context learning in LLM, and interestingly we find similar representation structures to Zhou et al.
-(2024) and Kantamneni & Tegmark (2025).
+Unlike prior work, we analyze addition in the ICL setting for LLMs, and interestingly we find similar representation structures to them.
 Interpreting in-context learning.
@@
 Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads.
-Motivated by this, we study function vector heads in details for a family of ICL tasks, introducing a novel optimization method, which achieves better performance than the method in Todd et al.
+Motivated by this, we study function vector heads in detail for a family of few-shot ICL tasks, introducing a novel optimization method, which achieves better performance than the method in Todd et al.
 (2024).
 Another difference from Todd et al.
-(2024) is that our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.
+(2024) is that our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information, which allows for a clearer understanding of ICL mechanism.
 Causal analysis.
@@
 7 Discussion
-In this work, we provided a detailed mechanistic analysis of in-context learning for addition tasks in Llama-3-8B.
+In this work, we provided a detailed mechanistic analysis of in-context learning for addition tasks.
 We found that a small number of attention heads operating in low-dimensional subspaces can extract, represent, and aggregate ICL task information in structured and interpretable ways.
@@
   4. 4.
-    Examine this subspace qualitatively, which uncovered periodic patterns in the activation space (§4.2) that decomposed into interpretable subspaces encoding unit-digit and magnitude information (§4.3).
+    Examine subspace qualitatively, which uncovered periodic patterns in activation space (§4.2) that decomposed into interpretable subspaces encoding unit-digit and magnitude information (§4.3).
   5. 5.
@@
 For the final step that involves a qualitative examination of the subspaces, future work could explore automating this step using AI systems.
-There are a few other questions stemming from our work.
-First, the three important heads are capable of individually representing and extracting information, but Llama3-8B relies on all of them—how does it jointly aggregate information from these three heads to produce a final answer?
-Second, we observe that heads implement a “self-correction” mechanism indicated by the negative correlation between examples signals, but the exact mechanism of this self-correction remains to be discovered.
-Our findings demonstrate how a focused study can reveal sophisticated latent structures and delicate computational strategies within a LLM.
-We hope this motivates more work to dive deep into the LLM inner structures and obtain more understanding of LLM latent spaces and computations.
 Acknowledgments
-We thank Lisa Dunlap, Alex Pan, Tony Lian, Anya Ji, Jiaxin Ge, Junyi Zhang, Hanlin Zhu and Shu Liu for their helpful feedback and suggestions.
+We thank Lisa Dunlap, Alex Pan, Tony Lian, Ryan Wang, Jiaxin Ge, Anya Ji, Junyi Zhang, Hanlin Zhu, Shu Liu and Baifeng Shi for their helpful feedback and suggestions.
 KY is supported by the Vitalik Buterin Ph.D.
@@
 JS is supported by Open Philanthropy, Simons Foundation, and NSF.
-Appendix A Additional table in §3.2
+Appendix A Supplement for §3.1
+A.1 Additional figure for §3.1
+In §3.1, we developed a global optimization method to identify $33$ significant heads for Llama-3-8B-instruct.
+Here we visualize our optimized coefficients as well as the previous method, also comparing their intervention accuracy in Figure 3.
+(a) Coefficients of heads.
+(b) Average indirect effects of heads (Todd et al., 2024).
+Figure 3: Comparison of significant heads identified by our optimized coefficients (left) and by average indirect effects (AIE) from the previous method (Todd et al., 2024) (right).
+Colors indicate the magnitude of each head’s importance (coefficients or AIE) on Llama-3-8B-instruct.
+The top $33$ heads identified by both methods are highlighted with frames ($13$ heads common across both methods in red and other $20$ heads in blue).
+Our identified heads yield an intervention accuracy of $0.85$, compared to the previous method’s accuracy of $0.31$.
+Both methods select heads from similar layers, but our optimization approach is significantly more effective.
+A.2 Additional models for §3.1
+We also train coefficients to get sets of important heads responsible for add-k on Llama-3.2-3B-instruct, Llama-3.2-3B, and Qwen-2.5-7B.
+In all cases, we achieve better accuracies with fewer heads than Todd et al.
+(2024).
+We report our results in Table 2.
+Model Llama-3.2-3B-instruct Llama-3.2-3B Qwen-2.5-7B
+ICL accuracy 0.50 0.64 0.91
+Our weighted-heads accuracy 0.66 0.83 0.61
+Our top-heads accuracy 0.62 0.75 0.34
+Todd et al. (2024)’s top-heads accuracy 0.20 0.1 0.12
+Table 2: Accuracies comparison between our method and Todd et al.
+(2024)’s method as well as the baseline ICL accuracy on add-k task cross other models.
+Our weighted-heads accuracy is the intervention accuracy achieved by the weighted sum of all heads where the weights are the raw coefficients at the last epoch of our training; our top-heads accuracy is the intervention accuracy achieved by the sum of top heads selected by coefficients at the last epoch of our training; and the Todd et al.
+(2024)’s top-heads accuracy is the intervention accuracy achieved by the sum of top heads selected by their average indirect effect.
+Here we choose the number of top heads as the one giving the highest accuracy for each case.
+For Llama-3 models, we need to choose around 30 heads while Todd et al.
+(2024)’s method needs to choose around 60 heads; for Qwen-2.5 model, we both need to choose $210$ heads.
+Our weighted-heads accuracy get significantly higher accuracy than Todd et al.
+(2024)’s, which indicates our method can find better function vectors than Todd et al.
+(2024)’s.
+Our top-heads accuracy are also higher with fewer heads than Todd et al.
+(2024)’s, which indicates our method can also find more effective set of heads for ICL tasks.
+Appendix B Supplement for §3.2
+B.1 Additional table for §3.2
 In §3.2, we did systematic ablation studies to narrow down the tasks to three main heads from $33$ significant heads.
 The first step of the ablation studies is layer-wise ablation, where we mean-ablate the significant heads in a subset of layers.
-We include the experimental results here, which narrow down to layer 13 and 15.
+We include the experimental results here, which narrow down to layer 13 and 15 in Table 3.
 Layer Accuracy
@@
 $\{14\}$ 0.03
-$[0,31]\setminus\{13,15\}$ 0.05
-Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers.
+$31]\setminus\{13,15\} 0.0$5
+Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers on Llama-3-8B-instruct.
 We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.87$, while other combinations lead to substantial drops in accuracy, especially when mean-ablating layers $13$ and $15$ (in blue).
-Appendix B Additional figures in §4
-In §4.2 and §4.3, we focused on head 1 for simplicity of exposition. Here we include the analogous experimental results for head 2 and 3.
-B.1 Additional figures for §4.2
-In §4.2, we found six trigonometric functions that can be linearly fitted by the coordinate functions of head 1 (Figure 6). Here we show analogous experimental results for head 2 and 3.
-Figure 10: The coordinate functions (inner product with feature directions) of head 2’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
-Figure 11: The coordinate functions (inner product with feature directions) of head 3’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
-B.2 Additional figures for §4.3
-In §4.3, we showed experimental results validating the hypothesis (ii) for head 1 (Figure 7).
-Here we show analogous experimental results validating the hypotheses (i) and (iii) for head 1, as well as all three hypotheses for heads 2 and 3.
-Recall that the three hypotheses are as follows:
+B.2 Additional models for §3.2
+We do the same ablation experiments for other three models, narrowing down to one layer and three heads for all of them.
+For two Llama-3.2 models, we narrow down to layer 14 and heads (14, 1), (14, 2), and (14, 12), the sum of which achieves accuracy $0.60$ and $0.70$.
+For Qwen-2.5-7B model, we narrow down to layer 21 and heads (21, 0), (21, 2), and (21, 5), the sum of which achieves accuracy $0.29$.
+Note that all of accuracies of these three heads are higher than the corresponding accuracies by Todd et al.
+(2024).
+We report the specific accuracies under different ablation setups in Tables 4, 5, 6, and 7.
+Layer Llama-3.2-3B-instruct Accuracy Llama-3.2-3B Accuracy
+No intervention 0.50 0.64
+$[0,27]$ 0.62 0.75
+$\{14\}$ 0.60 0.70
+Other single layer (max) 0.06 0.05
+Table 4: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers on Llama-3.2-3B-instruct and Llama-3.2-3B models.
+Both models narrow down to the layer $14$, which achieves significantly higher accuracy than any other single layer.
+Head Coefficients Llama-3.2-3B-instruct Accuracy Llama-3.2-3B Accuracy
+$(14,1)$ 5 / 5 0.78 0.95
+$(14,2)$ 6 / 4 0.51 0.61
+$(14,12)$ 4 / 4 0.26 0.26
+Other single head Optimal 0.03 0.08
+Table 5: Intervention accuracies for keeping the selected heads scaled by the corresponding coefficients and mean-ablating all the other significant heads on Llama-3.2-3B-instruct and Llama-3.2-3B models.
+We narrow down them both to the same three heads, which achieve significantly higher accuracy than any other single head in layer $14$.
+Layer Qwen-2.5-7B Accuracy
+No intervention 0.91
+$[0,27]$ 0.34
+$\{21\}$ 0.29
+Other single layer (max) 0.03
+Table 6: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers on Llama-3.2-3B-instruct and Llama-3.2-3B models.
+Both models narrow down to the layer $14$, which achieves significantly higher accuracy than any other single layer.
+Head Coefficient Qwen-2.5-7B Accuracy
+$(21,5)$ 4 0.29
+$(21,0)$ 3 0.18
+$(21,2)$ 4 0.15
+Other single head Optimal 0.06
+Table 7: Intervention accuracies for keeping the selected heads scaled by the corresponding coefficients and mean-ablating all the other significant heads on Qwen-2.5-7B.
+We narrow down to three heads, which achieve significantly higher accuracy than any other single head in layer 2121.
+Appendix C Supplement for §4.1
+We perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance on Llama-3-8B-instruct (Figure 4(a)), and similarly for the three models (Figures 4(b), 4(c) and 4(d)).
+(a) Llama-3-8B-instruct
+(b) Llama-3.2-3B-instruct
+(c) Llama-3.2-3B
+(d) Qwen-2.5-7B
+Figure 4: Explained variance ratio vs.
+number of PCs for each head across models.
+The first six PCs make up most of the explained variance (97%) for Llama-3 models, and the first eight PCs do so for Qwen-2.5-7B.
+Appendix D Supplement for §4.2
+We first present additional figures for Llama-3-8B-Instruct, followed by results for the other three models.
+The main finding is that three heads in each Llama-3.2 model and one head in the Qwen-2.5 model consistently encode periodic patterns through trigonometric functions.
+By contrast, the remaining two heads in Qwen-2.5 do not encode periodicity; instead, their coordinate functions exhibit distinct behaviors across the intervals $[1,10]$, $[11,20]$, and $[21,30]$.
+We conjecture that this difference arises from tokenization: Qwen encodes numbers digit by digit, leading to discontinuities across each 10-interval, whereas the Llama family represents every number below 100 as a single token.
+D.1 Additional figures for §4.2
+In §4.2, we found six trigonometric functions that can be linearly fitted by the coordinate functions.
+Here we first supplement the plot from which we observe the periodic pattern of the coordinate functions for the three heads (Figure 5) and then show the fitting functions for three heads (Figure 6).
+(a) Head (15, 2)
+(b) Head (15, 1)
+(c) Head (13, 6)
+Figure 5: Coordinates of three heads’ vectors (inner products with PCs) for the first six PCs across different add-$k$ tasks on Llama-3-8B-instruct.
+Periodic patterns are visible in the first few PCs of each head.
+(a) Head (15, 2)
+(b) Head (15, 1)
+(c) Head (13, 6)
+Figure 6: Coordinate functions of three heads can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ on Llama-3-8B-instruct.
+D.2 Llama-3.2-3B-instruct for §4.2
+The Llama family of models show similar results. We show results for Llama-3.2-3B-instruct here and omit Llama-3.2-3B.
+(a) Head (14, 1)
+(b) Head (14, 2)
+(c) Head (14, 12)
+Figure 7: Coordinates of three heads’ vectors (inner products with PCs) for the first six PCs across different add-$k$ tasks on Llama-3.2-3B-instruct. The first six PCs reveal clear periodic patterns.
+(a) Head (14, 1)
+(b) Head (14, 2)
+(c) Head (14, 12)
+Figure 8: Coordinate functions of three heads can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ on Llama-3.2-3B-instruct.
+D.3 Qwen-2.5-7B for §4.2
+One head (21, 0) shows similar periodic patterns while the other two heads do not encode periodicity; instead, their coordinate functions exhibit distinct behaviors across the intervals $[1,10]$, $[11,20]$, and $[21,30]$.
+We conjecture that this difference arises from tokenization: Qwen encodes numbers digit by digit, leading to discontinuities across each 10-interval, whereas the Llama family represents every number below 100 as a single token.
+(a) Head (21, 0)
+(b) Head (21, 2)
+(c) Head (21, 5)
+Figure 9: Coordinates of three heads’ vectors (inner products with PCs) for the first six PCs across different add-$k$ tasks on Qwen-2.5-7B.
+Head (21, 0) reveals periodic patterns, whereas Heads (21, 2) and (21, 5) show discontinuous behaviors across the intervals $[1,10]$, $[11,20]$, and $[21,30]$.
+(a) Head (21, 0)
+(b) Head (21, 2)
+(c) Head (21, 5)
+Figure 10: Six trigonometric functions with periods $2,5,10,10,25$, and $50$ fitted by coordinate functions of three heads on Qwen-2.5-7B.
+Only head (21, 0) fits the periodic functions well, while heads (21, 2) and (21, 5) do not, consistent with their non-periodic coordinate behaviors.
+Appendix E Supplement for §4.3
+Recall that the three hypotheses in §4.3 are as follows:
   1. (i)
@@
   3. (iii)
-    the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
-We first show the experimental results validating hypotheses (i) and (iii) for head 1.
-Projecting out of the parity direction doesn’t lead to high errors for the parity and the final answer across all tasks, which might be because parity is relatively easy to obtain (e.g., random choice leads to 0.5 accuracy).
-(a) Head 1 projected onto parity direction
-(b) Head 1 projected out of parity direction
-Figure 12: Validation of hypothesis (i) for head 1.
-(a) Head 1 projected onto magnitude subspace
-(b) Head 1 projected out of magnitude subspace
-Figure 13: Validation of hypothesis (iii) for head 1.
-We then show analogous results for head 2.
-(a) Head 2 projected onto parity direction
-(b) Head 2 projected out of parity direction
-Figure 14: Validation of hypothesis (i) for head 2.
-(a) Head 2 projected onto unit subspace
-(b) Head 2 projected out of unit subspace
-Figure 15: Validation of hypothesis (ii) for head 2.
-(a) Head 2 projected onto magnitude subspace
-(b) Head 2 projected out of magnitude subspace
-Figure 16: Validation of hypothesis (iii) for head 2.
-Finally, we show analogous results for head 3.
-The evidence for head 3 is slightly weaker than the first two heads, which aligns with the fact that the intervened accuracy of head 3 is slightly weaker than the ones of the first two heads (Tabel 1).
-(a) Head 3 projected onto parity direction
-(b) Head 3 projected out of parity direction
-Figure 17: Validation of hypothesis (i) for head 3.
-(a) Head 3 projected onto unit subspace
-(b) Head 3 projected out of unit subspace
-Figure 18: Validation of hypothesis (ii) for head 3.
-(a) Head 3 projected onto magnitude subspace
-(b) Head 3 projected out of magnitude subspace
-Figure 19: Validation of hypothesis (iii) for head 3.
+    the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
+We first show the experimental results validating hypothesis (ii) for head 1 (Figure 11), and then show analogous results for hypotheses (i) and (iii).
+Projecting out of the parity direction does not substantially increase errors for either the parity or the final answer across all tasks, which might be because parity is relatively easy to obtain (e.g., random choice leads to 0.5 accuracy).
+Figure 11: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto (left) and out of (right) the “unit subspace”.
+Projecting onto the unit subspace results in a low unit-digit error rate even when the final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully account for the final-answer errors.
+This confirms that the unit subspace specifically encodes the unit-digit signal.
+(a) onto parity direction
+(b) out of parity direction
+(c) onto magnitude subspace
+(d) out of magnitude subspace
+Figure 12: Validation of hypotheses (i) and (iii) for head 1.
+Each row shows results for the parity and magnitude subspaces (left: projection onto; right: projection out of).
+The unit-digit hypothesis (ii) is omitted here since it is already shown in Figure 11.
+(a) Head 2 onto parity direction
+(b) Head 2 out of parity direction
+(c) Head 2 onto unit subspace
+(d) Head 2 out of unit subspace
+(e) Head 2 onto magnitude subspace
+(f) Head 2 out of magnitude subspace
+Figure 13: Validation of hypotheses (i)–(iii) for head 2. Each row shows the projection effects for the parity, unit, and magnitude subspaces respectively.
+(a) Head 3 onto parity direction
+(b) Head 3 out of parity direction
+(c) Head 3 onto unit subspace
+(d) Head 3 out of unit subspace
+(e) Head 3 onto magnitude subspace
+(f) Head 3 out of magnitude subspace
+Figure 14: Validation of hypotheses (i)–(iii) for head 3. The evidence is slightly weaker than for heads 1 and 2, consistent with head 3’s lower intervened accuracy (Table 1).
+Appendix F Supplement for §5.2
+We show the strength and direction of the signals extracted from each individual token through a random mixed-k ICL prompt as an example below.
+Since we find out the Llama family of models behave similarly in the previous sections, we here just show results for Llama-3-8B-instruct and Qwen-2.5-7B as examples.
+F.1 Additional figures for §5.2
+All three heads for Llama-3-8B-instruct behave similarly in the signal strengths and directions. They all peak at the $y$ tokens and extract signal corresponding to $y_{i}-x_{i}$.
+(a) Head (15, 2)
+(b) Head (15, 1)
+(c) Head (13, 6)
+Figure 15: Llama-3-8B-instruct.
+For each head, the top panel shows the strength of task-signal contribution of each previous token $t$ to the final token, $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$.
+Decomposing it into two parts: (1) the norm of the extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$ and (2) the attention score from the final token $\alpha_{t}$, both consistently peak at the tokens $t=y_{i}$ (in bold).
+The bottom panel shows the inner product between the projected signal from each $y_{i}$ and the head vector $\tilde{h}_{k}$ (normalized to unit norm), which peaks at $k=y_{i}-x_{i}$, indicating that each head extracts $y_{i}-x_{i}$ from its corresponding demonstration $x_{i}\!\to\!y_{i}$.
+F.2 Qwen-2.5-7B for §5.2
+Notice that Qwen-2.5-7B has a different tokenizer from the Llama models: it tokenizes numbers digit by digit.
+The three heads behave qualitatively similarly—peaking at $y$ tokens and extracting $y_{i}-x_{i}$—though sometimes with an offset of $10$ or $20$.
+(a) Head (21, 0)
+(b) Head (21, 2)
+(c) Head (21, 5)
+Figure 16: Qwen-2.5-7B.
+For each head, the top panel shows $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$, decomposed into two parts: (1) the norm of extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$, and (2) the attention score from the final token $\alpha_{t}$.
+Both consistently peak at the tokens $t=y_{i}$ (in bold).
+The bottom panel shows the inner products between the projected signals from $y_{i}$ and head vectors $\tilde{h}_{k}$, which peak near $k=y_{i}-x_{i}$ (sometimes offset by 10 or 20), consistent with Qwen’s digit-wise tokenization.
+Appendix G Supplement for §5.3
+We report the concrete numeric variables computed in §5.3 for Llama-3-8B-instruct, Llama-3.2-3B, and Qwen-2.5-7B.
+The self-correction mechanism significantly exists in Llama family of models while partially exists in Qwen model.
+G.1 Additional table for §5.3
+The signals from any two demonstrations are mostly negatively correlated for all three heads, suggesting a self-correction mechanism on Llama-3-8B-instruct.
+Stat Neg Pos
+Avg -2.01 0.27
+Min -1.40 0.07
+Max -2.34 0.54
+(a) Head (15, 2)
+Stat Neg Pos
+Avg -1.95 0.05
+Min -1.65 0.00
+Max -2.17 0.20
+(b) Head (15, 1)
+Stat Neg Pos
+Avg -0.76 0.28
+Min -0.15 0.00
+Max -1.68 1.03
+(c) Head (13, 6)
+Table 8: Llama-3-8B-instruct: Statistics (min, max, average) of the absolute values of negative and positive correlation sums over the 30 tasks for three heads.
+The negative correlation sum is significantly higher for all heads, indicating that signals from any two demonstrations are mostly negatively correlated — a hallmark of the self-correction mechanism.
+G.2 Llama-3.2-3B for §5.3
+The signals from any two demonstrations are mostly negatively correlated for the two stronger heads, consistent with the self-correction mechanism observed in Llama-3-8B-instruct.
+Stat Neg Pos
+Avg -2.12 0.23
+Min -1.96 0.00
+Max -2.41 0.53
+(a) Head (14, 1)
+Stat Neg Pos
+Avg -2.07 0.05
+Min -1.89 0.00
+Max -2.24 0.22
+(b) Head (14, 2)
+Stat Neg Pos
+Avg 0.27 0.75
+Min 0.00 0.00
+Max -0.90 0.75
+(c) Head (14, 12)
+Table 9: Llama-3.2-3B: Statistics (min, max, average) of the absolute values of negative and positive correlation sums over the 30 tasks for three heads.
+Heads (14, 1) and (14, 2) show clear negative correlation dominance, while head (14, 12) does not — consistent with its weaker task accuracy (Table 5).
+G.3 Qwen-2.5-7B for §5.3
+For Qwen-2.5-7B, the correlation pattern varies by head.
+The head encoding periodic patterns (21, 0) still shows strong negative correlations, while the others exhibit weaker or even positive correlations—indicating that the self-correction mechanism only partially exists.
+Stat Neg Pos
+Avg -1.84 0.08
+Min -1.44 0.00
+Max -2.15 0.67
+(a) Head (21, 0)
+Stat Neg Pos
+Avg -0.64 0.50
+Min -1.44 0.03
+Max -2.19 1.57
+(b) Head (21, 5)
+Stat Neg Pos
+Avg -0.37 2.62
+Min 0.00 0.00
+Max -1.85 5.54
+(c) Head (21, 2)
+Table 10: Qwen-2.5-7B: Statistics (min, max, average) of the absolute values of negative and positive correlation sums over the 30 tasks for three heads.
+Head (21, 0) shows clear negative correlations, while Heads (21, 5) and (21, 2) show mixed or positive correlations, suggesting that self-correction partially exists in Qwen-2.5-7B.